# Personalize Your Gaussian: Consistent 3D Scene Personalization from a Single Image

### v1

@@ Abstract
Personalizing 3D scenes from a single reference image enables intuitive user-guided editing, which requires achieving both multi-view consistency across perspectives and referential consistency with the input image. However, these goals are particularly challenging due to the viewpoint bias caused by the limited perspective provided in a single image. Lacking the mechanisms to effectively expand reference information beyond the original view, existing methods of image-conditioned 3DGS personalization often suffer from this viewpoint bias and struggle to produce consistent results. Therefore, in this paper, we present Consistent Personalization for 3D Gaussian Splatting (CP-GS), a framework that progressively propagates the single-view reference appearance to novel perspectives. In particular, CP-GS integrates pre-trained image-to-3D generation and iterative LoRA fine-tuning to extract and extend the reference appearance, and finally produces faithful multi-view guidance images and the personalized 3DGS outputs through a view-consistent generation process guided by geometric cues. Extensive experiments on real-world scenes show that our CP-GS effectively mitigates the viewpoint bias, achieving high-quality personalization that significantly outperforms existing methods. The code will be released at https://github.com/Yuxuan-W/CP-GS.
@@ 1 Introduction
  • To mitigating the viewpoint bias, we propose a coarse-to-fine appearance propagation framework that progressively expands the single-view reference appearance to novel perspectives, generating guidance images with faithful referential consistency and strong multi-view consistency.
@@ 3.1 Problem Definition
where $\mathcal{R}$ denotes the rendering function [13]. This paradigm requires the multi-view guidance images $\mathcal{I}^{\text{gde}}$ to satisfy two key properties: multi-view consistency across $\mathcal{V}$ to prevent optimization conflicts, and referential appearance consistency to the $\mathcal{I}^{\text{ref}}$ to fulfill the personalization objective. Our CP-GS is designed to explicitly ensures both consistency to achieve high-quality 3D personalization.
@@ 3.2 Coarse Guidance Generation
As noted in Sec. 1, the limited perspective of a single reference image fails to provide sufficient geometric and coherent appearance information for constructing a consistent multi-view representation. Therefore, in the first stage, we leverage an off-the-shelf image-to-3D generation model TRELLIS [19], pre-trained on large-scale CGI-style 3D datasets [20, 21], to produce a coarse guidance scene that expands the reference into a rough yet multi-view consistent representation. As illustrated in the top-left of Figure 3, the reference image is fed into the pre-trained TRELLIS to generate the corresponding 3D rough asset, which is then integrated into the source scene to replace or augment the user-specific target region. Here we provide two integration modes: (1) Adding new content – the user provides a 3D bounding box specifying the object’s position and scale; (2) Replacing existing content – the target bounding box is extract from the existing content via PCA [66], and the generated asset is fitted accordingly. This coarse guidance provides a plausible 3D geometry and establishing a rough yet view-consistent appearance that serves as a structural prior for subsequent stages.
@@ 3.4 View-consistent Generation
In the final stage, we adopt a view-consistent generation straregy based on a pre-trained Flow Transformer [24], combining the coarse guidance (Sec. 3.2) and the iteratively trained LoRA module (Sec. 3.3) to produce the final consistent guidance images. As shown in Figure 3, we begin by rendering the coarse guidance into multi-view images and converting them into noisy latents using rectified-flow inversion [25] to encode both the appearance and geometry. Serving as the starting point for subsequent generation process, these latents are then fed into the Flow Transformer along with rendered depth maps, which provide geometric cues to align appearance generation with the underlying structure and enhance multi-view consistency.
Inspired by [6, 68], we introduce an epipolar-constrained token replacement mechanism, to promote multi-view consistency by unifying foreground tokens across views that correspond to the same 3D locations. We perform token replacement in the early dual-stream blocks of the Flow Transformer, where visual tokens are explicitly maintained and can be directly modified. During generation, we automatically select a set of key views with minimal overlap to ensure full coverage, and extract foreground pixel indices in all viewpoints using masks from the coarse guidance. As illustrated in Figure 4(b), for each foreground token in non-key views, we compute its epipolar lines on the two nearest key views and replace it with an interpolation of the most similar tokens along those epipolar lines, weighted by the camera distance to each key view. Given a non-key frame, the interpolated token $\mathbf{f}^{\prime}(\mathbf{u})$ at pixel $\mathbf{u}$ used to replace the original token $\mathbf{f}(\mathbf{u})$ is computed using the foreground tokens of the two nearest key frames $\mathbf{k}_{i}$, indexed by $i\in\{1,2\}$. Letting $c$ denote the non-key camera and $l_{\mathbf{u}\rightarrow i}$ denote the epipolar line of $\mathbf{u}$ in each key view, the token $\mathbf{f}\textquoteright(\mathbf{u})$ is computed as:
@@
where $\mathcal{D}(c,c_{i})$ represents the camera distance from $c$ to each key view’s camera $c_{i}$. This mechanism effectively alleviates cross-view variance, producing fine-grained guidance images with multi-view consistency and faithful reference alignment. These images then supervise the 3DGS parameter updating of the coarse guidance, yielding the final personalized 3DGS result of our CP-GS framework.

### v3

@@ Abstract
Personalizing 3D scenes from a single reference image enables intuitive user-guided editing, which requires achieving both multi-view consistency across perspectives and referential consistency with the input image. However, these goals are particularly challenging due to the viewpoint bias caused by the limited perspective provided in a single image. Lacking the mechanisms to effectively expand reference information beyond the original view, existing methods of image-conditioned 3DGS personalization often suffer from this viewpoint bias and struggle to produce consistent results. Therefore, in this paper, we present Consistent Personalization for 3D Gaussian Splatting (CP-GS), a framework that progressively propagates the single-view reference appearance to novel perspectives. In particular, CP-GS integrates pre-trained image-to-3D generation and iterative LoRA fine-tuning to extract and extend the reference appearance, and finally produces faithful multi-view guidance images and the personalized 3DGS outputs through a view-consistent generation process guided by geometric cues. Extensive experiments on real-world scenes show that our CP-GS effectively mitigates the viewpoint bias, achieving high-quality personalization that significantly outperforms existing methods.
@@ 1 Introduction
  • To mitigate the viewpoint bias, we propose a coarse-to-fine appearance propagation framework that progressively expands the single-view reference appearance to novel perspectives, generating guidance images with faithful referential consistency and strong multi-view consistency.
@@ 3.1 Problem Definition
where $\mathcal{R}$ denotes the rendering function [13]. This paradigm requires the multi-view guidance images $\mathcal{I}^{\text{gde}}$ to satisfy two key properties: multi-view consistency across $\mathcal{V}$ to prevent optimization conflicts, and referential appearance consistency to the $\mathcal{I}^{\text{ref}}$ to fulfill the personalization objective. Our CP-GS is designed to explicitly ensure both consistency to achieve high-quality 3D personalization.
@@ 3.2 Coarse Guidance Generation
As noted in Sec. 1, the limited perspective of a single reference image fails to provide sufficient geometric and coherent appearance information for constructing a consistent multi-view representation. Therefore, in the first stage, we leverage an off-the-shelf image-to-3D generation model TRELLIS [19], pre-trained on large-scale CGI-style 3D datasets [20, 21], to produce a coarse guidance scene that expands the reference into a rough yet multi-view consistent representation. As illustrated in the top-left of Figure 3, the reference image is fed into the pre-trained TRELLIS to generate the corresponding 3D rough asset, which is then integrated into the source scene to replace or augment the user-specific target region. We provide two integration modes: (1) Adding new content – the user provides a 3D bounding box specifying the object’s position and scale; (2) Replacing existing content – the target bounding box is extracted from the existing content via PCA [66], and the generated asset is fitted accordingly. This coarse guidance provides a plausible 3D geometry and establishing a rough yet view-consistent appearance that serves as a structural prior for subsequent stages.
@@ 3.4 View-consistent Generation
In the final stage, we adopt a view-consistent generation strategy based on a pre-trained Flow Transformer [24], combining the coarse guidance (Sec. 3.2) and the iteratively trained LoRA module (Sec. 3.3) to produce the final consistent guidance images. As shown in Figure 3, we begin by rendering the coarse guidance into multi-view images and converting them into noisy latents using rectified-flow inversion [25] to encode both the appearance and geometry. Serving as the starting point for subsequent generation process, these latents are then fed into the Flow Transformer along with rendered depth maps, which provide geometric cues to align appearance generation with the underlying structure and enhance multi-view consistency.
Inspired by [6, 68], we introduce an epipolar-constrained token replacement mechanism, to promote multi-view consistency by unifying foreground tokens across views that correspond to the same 3D locations. We perform token replacement in the early dual-stream blocks of the Flow Transformer, where visual tokens are explicitly maintained and can be directly modified. During generation, we automatically select a set of key views with minimal overlap to ensure full coverage, and extract foreground pixel indices in all viewpoints using masks from the coarse guidance. As illustrated in Figure 4(b), for each foreground token in non-key views, we compute its epipolar lines on the two nearest key views and replace it with an interpolation of the most similar tokens along those epipolar lines, weighted by the camera distance to each key view. Given a non-key frame, the interpolated token $\mathbf{f}^{\prime}(\mathbf{u})$ at pixel $\mathbf{u}$ used to replace the original token $\mathbf{f}(\mathbf{u})$ is computed using the foreground tokens of the two nearest key frames $\mathbf{k}_{i}$, indexed by $i\in\{1,2\}$. Letting $c$ denote the non-key camera and $l_{\mathbf{u}\rightarrow i}$ denote the epipolar line of $\mathbf{u}$ in each key view, the token $\mathbf{f}’(\mathbf{u})$ is computed as:
@@
where $\mathcal{D}(c,c_{i})$ represents the camera distance from $c$ to each key view’s camera $c_{i}$. This mechanism effectively alleviates cross-view variance, producing guidance images with strong multi-view consistency and faithful reference alignment. These images then supervise the 3DGS parameter updating of the coarse guidance, yielding the final personalized 3DGS result of our CP-GS framework.
@@ 5 Conclusion
Appendix A More Implementation Details
A.1 Step-by-step Demonstration of CP-GS
In our main paper, we introduce our CP-GS personalization framework that progressively propagate the single-view reference image appearance to novel perspectives. In Algorithm 1, we provide a step-by-step demonstration of our entire pipeline for content adding or replacement, including the coarse guidance generation, the iterative LoRA fine-tuning, the view-consistent generation of guidance images, and the final 3DGS optimization that produces our personalized 3DGS results.
Algorithm 1 Step-by-Step Pipeline of CP-GS
1:Input: Source 3DGS scene $\mathcal{G}^{\text{src}}$ with view set $\mathcal{V}$, single-view reference image $\mathcal{I}^{\text{ref}}$, and user-specific target region.
2:Forward $\mathcal{I}^{\text{ref}}$ into the pre-trained image-to-3D [19] model, generating a coarse 3D asset $\mathcal{G}^{\text{coar}}$.
3:Integrate $\mathcal{G}^{\text{coar}}$ into the target region of $\mathcal{G}^{\text{src}}$, producing coarse guidance scene $\mathcal{G}$.
4:Render $\mathcal{G}$ from $\mathcal{V}$ into multi-view images $\mathcal{I}^{\text{rend}}$, depth maps $\mathcal{D}^{\text{rend}}$, and masks $\mathcal{M}^{\text{rend}}$ indicating $\mathcal{G}^{\text{coar}}\in\mathcal{G}$.
5:Initialize the image model $\mathcal{F}$ by a pre-trained Flux [24] model and LoRA module.
6:Initialize the training image set for LoRA fine-tuning as $\mathcal{I}^{\text{train}}_{1}=\{\mathcal{I}^{\text{ref}}\}$.
7:for t=1,2,Tt=1,2,...T do ($T=2$ as default):
8:        Fine-tune the LoRA module in $\mathcal{F}$ using $\mathcal{I}^{\text{train}}_{t}$, where the background of $\mathcal{I}^{\text{train}}_{t}$ is excluded.
9:        Translate the multi-view $\mathcal{I}^{\text{rend}}$ into $\mathcal{I}^{\text{trans}}$ by $\mathcal{I}^{\text{trans}}=\mathcal{F}(\mathcal{I}^{\text{rend}})$.
10:      For $I^{\text{trans}}_{i}\in\mathcal{I}^{\text{trans}}_{t}$, compute similarity $\mathbf{S}_{i}=\underset{j}{\sum}\>\mathbf{S}_{\text{RoMa}}(I^{\text{trans}}_{i},I^{\text{train}}_{j})$ towards $\mathcal{I}^{\text{train}}_{t}=\{I^{\text{train}}_{j}\}$.
11:      Extend $\mathcal{I}^{\text{train}}_{t+1}=\mathcal{I}^{\text{train}}_{t}\cup\{\underset{i}{\operatorname*{arg\,min}}\>\mathbf{S}_{i}\}$ by the one translation with minimal $\mathbf{S}_{i}$.
12:Select multiple key views $\mathbf{k}$ from $\mathcal{I}^{\text{rend}}$ by their perspective coverage.
13:For the rest non-key view $\mathbf{f}$, find the 22 nearest key views $\mathbf{k}_{i},i\in\{1,2\}$ by camera distance $\mathcal{D}(c,c_{i})$.
14:Compute the epipolar lines $\l _{\mathbf{u}\rightarrow i}$ in $\mathbf{k}_{i}$ corresponding to each pixel $\mathbf{u}$ in non-key views $\mathbf{f}$.
15:Convert $\mathcal{I}^{\text{rend}}$ to noisy latents $\mathcal{Z}^{\text{rend}}$ using Rectified Flow Inversion [25].
16:Forward $\mathcal{Z}^{\text{rend}}$, $\mathcal{D}^{\text{rend}}$, and $\mathcal{M}^{\text{rend}}$ into $\mathcal{F}$, generating guidance images $\mathcal{I}^{\text{gde}}=\mathcal{F}(\mathcal{Z}^{\text{rend}},\mathcal{D}^{\text{rend}},\mathcal{M}^{\text{rend}})$.
17:      In dual-stream blocks of $\mathcal{F}$, for pixel $\mathbf{u}\in\mathcal{M}^{\text{rend}}$ do:
18:           $\mathbf{f}(\mathbf{u})\leftarrow\sum_{i}\mathbf{k}_{i}(\mathbf{v}_{i})\mathcal{D}(c,c_{i})/\sum_{i}\mathcal{D}(c,c_{i}),\>\>\text{where}\>\>\mathbf{v}_{i}=\underset{\mathbf{v}\in l_{\mathbf{u}\rightarrow i}}{\operatorname*{arg\,max}}\>\langle\mathbf{f}(\mathbf{u}),\mathbf{k}_{i}(\mathbf{v})\rangle$.
19:Optimize the coarse guidance: $\mathcal{G}^{\text{edit}}=\>\underset{\mathcal{G}}{\mathrm{argmin}}\>\sum_{v\in\mathcal{V}}{(\lambda_{1}\mathcal{\mathcal{L}_{\text{MAE}}}(\mathcal{R}(\mathcal{G},v),\>\mathcal{I}^{\text{gde}})+\lambda_{2}\mathcal{\mathcal{L}_{\text{LPIPS}}}(\mathcal{R}(\mathcal{G},v),\>\mathcal{I}^{\text{gde}}))}$.
20:Output: personalized 3DGS scene $\mathcal{G}^{\text{edit}}$.
A.2 Coarse Asset Integration
In our main paper, we introduce two integration strategies for generating coarse guidance and support three types of personalization from a single reference image.
For adding a new object to the scene (e.g., the man wearing sunglasses), we allow the user to specify a 3D bounding box indicating the desired location and scale of the inserted object. The coarse asset generated by the image-to-3D model [19] is then placed accordingly. For style transfer, coarse guidance is not required since there are no significant geometric changes between the source and personalization results. Instead, we directly extract the reference appearance from the input image and perform the view-consistent generation stage to produce the final guidance images.
For replacing an existing object (e.g., replacing the bear with a dog) in the source scene, we first remove the original content by partially adopting the deletion pipeline from GaussianEditor [1]. Specifically, we use SegmentAnything [75] to segment the foreground from multi-view renderings, then project the segmented regions into 3D space to identify the corresponding Gaussians. These Gaussians are subsequently pruned from the scene. Then we use SDXL [15] to inpaint and fix the resulting hole caused by the deletion, following the repairing pipeline of GaussianEditor [1].
To insert the new object, we provide a PCA-based [66] bounding box extraction tool to assist in estimating the bounding box of the original object (e.g., the bear). We first determine the up direction from the “ground” Gaussians extracted using the same segmentation method, identifying it as the principal component axis corresponding to the smallest eigenvalue in PCA. The forward direction is then derived from the removed original object by finding the PCA axis with the largest eigenvalue within the plane orthogonal to the up direction. The right direction is computed as the cross product of the forward and up vectors. We project all Gaussians of the original object onto these axes and compute the min–max range along each axis to define the 3D bounding box, discarding outliers by retaining the central 98% of the data along each dimension. This PCA-based bounding box extraction utility works well in most cases where the inserted object is expected to follow the original orientation and is placed upright. For custom insertions, users can manually adjust the estimated bounding box to indicate the desired position and alignment.
A.3 Extension of LoRA Training Set
For the translation process in our iterative LoRA [67] fine-tuning, we use the same positive prompt as in the LoRA training while adopting a unified negative prompt: “ugly, deformed, disfigured, poor details, bad anatomy, cartoon, CGI, unrealistic”, to suppress undesired artifacts and improve translation quality. In the selective augmentation of the LoRA training set, to compute the RoMa [23] similarity score between each translated image and each training image, we compute the average confidence of fixed 10,000 matching points predicted by the RoMa model, following the default configuration defined in their official implementation.
A.4 Details of User Study
In this section, we provide additional details regarding our User Study. As illustrated in Figure 9, the study consists of two types of questions aimed at evaluating two key aspects of each personalized result: visual quality and referential alignment with the input image. To ensure fairness and avoid bias, the method names are anonymized and the order of answer choices is randomized for each question. Participants are asked to indicate whether they are satisfied with each presented result. For every method, we aggregate the responses and compute the average satisfaction rate across all samples and participants. The study comprises 20 questions for each question type, spanning 20 representative samples and 60 personalized results from three compared methods. In total, we collected responses from 33 participants with diverse academic and professional backgrounds. This carefully designed questionnaire, combined with a broad participant base, ensures the reliability of the results.
Figure 9: The interface of our user study includes two types of questions: (1) Given the rendered source scene and the 3DGS results from each compared method, users are asked whether they are satisfied with the visual quality of each result. (2) Given the reference image and the renderings of each 3DGS result, users are asked to provide their satisfactory on the referential alignment.
Appendix B Extensive Experiments
B.1 More Qualitative Results
As a supplement to our main paper, we present additional personalization results produced by CP-GS in Figure 10, where CP-GS effectively mitigates viewpoint bias and delivers high-quality personalization outcomes for each sample. These results further demonstrate the adaptability of CP-GS across diverse and complex scenarios with varied reference images.
Table 2: Quantitative ablation study of our CP-GS across four configurations: the Coarse Guidance directly integrating the generated asset, the Single-view LoRA trained on the single reference image, the Iterative LoRA with iterative LoRA fine-tuning, and the Full Version further incorporating the epipolar-constrain token replacement mechanism.
Methods Coarse Guidance Single-view LoRA Iterative LoRA Full Version
[…]
B.2 Ablated Quantitative Evaluations
In our main paper, we conducted a qualitative ablation study on each component of CP-GS, including the coarse guidance, iterative LoRA fine-tuning, and epipolar-constrained token replacement. Here, we extend this analysis with a quantitative evaluation using the same test samples. Table 2 compares four configurations: (1) Coarse Guidance, which directly integrates the image-to-3D asset generated by TRELLIS [19]; (2) Single-view LoRA, which trains the LoRA module solely on the single reference image for appearance refinement; (3) Iterative LoRA, which applies our iterative LoRA fine-tuning but excludes the epipolar-constrained token replacement; and (4) the Full Version CP-GS, which incorporates all components. We adopt the same evaluation metrics as in the main paper: CLIP [73] and DINO [72] image-to-image similarity, as well as CLIP directional similarity [74]. Lower values across these metrics indicate better alignment between the personalization outputs and the reference image, reflecting stronger referential consistency that fulfills the objective of the task.
As shown in Table 2, the comparison of quantitative results align with our qualitative findings, confirming that the Full Version of CP-GS outperforms all ablated variants. Notably, the iterative LoRA fine-tuning contributes most significantly to performance gains by effectively propagating fine-grained reference appearance to novel views (see Iterative LoRA vs. Single-view LoRA). The addition of epipolar-constrained token replacement further enhances multi-view consistency and improves the visual quality of the final 3DGS output (see Full Version vs. Iterative LoRA). An interesting observation is that the Single-view LoRA performs noticeably worse than the Coarse Guidance, suggesting that naive refinement using a viewpoint-biased image generation model can degrade the quality of the coarse 3DGS asset generated from image-to-3D model [19]. In contrast, both the Iterative LoRA and Full Version show marked improvements over the Coarse Guidance, demonstrating their effectiveness in addressing the viewpoint bias caused by limited reference perspective and generating consistent editing guidance. These findings confirm the importance of progressive reference appearance propagation for achieving strong referential alignment in final 3DGS personalization results.
Appendix C Other Discussions
C.1 Limitation Discussion
While CP-GS demonstrates compelling performance in 3DGS personalization conditioned on single-view reference image, several limitations remain. First, the iterative LoRA fine-tuning, though efficient, still requires multiple inference and training rounds, which hinders the application in large-scale batch editing scenarios. Besides, faithfully reproducing extremely intricate visual details, including fine-grained patterns and embedded text, remains a challenging aspect for the current LoRA module. Second, due to the lack of effective automatic 3DGS [13] insertion method, our method still relies on user-provided bounding boxes in part of scenarios. In future work, we plan to develop automatic insertion strategies and enhance scalability by distilling the iterative process into a single forward pass.
C.2 Potential Societal Impacts
Our CP-GS framework offers several positive societal implications. By enabling high-quality 3D scene personalization from only a single reference image, it significantly reduces the dependency on extensive image collections or manual 3D modeling, thereby lowering both the cost and workload typically required from human artists. This contributes to a more accessible and efficient content creation process, aligning with the goals of sustainable and green AI.
On the other hand, the ability to easily personalize 3D content may raise concerns regarding potential misuse, such as the generation of inappropriate or harmful scenes involving graphic, violent, or NSFW elements. To mitigate this risk, CP-GS builds upon the diffuser [69] codebase that inherits safety mechanisms—such as NSFW content filters—from the underlying model [24]. These filters help ensure that the personalized outputs remain within acceptable and responsible usage boundaries.
Figure 10: Supplementary personalization result of our CP-GS, demonstrating high-quality 3DGS scene customization that faithfully align with the reference image across various scenarios.